OpenStudy (anonymous):

7b+3-4b+3-3(b+4) How do you get your answer?

OpenStudy (michele_laino):

please aply the ditributive property in order to simpliy the 5-th term, namely: \[ - 3(b + 4) = ...?\]

OpenStudy (michele_laino):

oops...please apply...

OpenStudy (anonymous):

\[7b+3-4b=3-3b-12\] Okay, Grammar is often ignored during social conversations.

OpenStudy (michele_laino):

I got this: \[\begin{gathered} 7b + 3 - 4b + 3 - 3\left( {b + 4} \right) = \hfill \\ = 3b + 6 - 3b - 12 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

I'm sorry I wrote it wrong the first time, its suppose to be \[7b+3-4b=3-3(b+4)\]

OpenStudy (michele_laino):

ok! then you are right!

OpenStudy (anonymous):

How? My teacher told us to use FOIL for 3-3(b+4) Should I?

OpenStudy (anonymous):

FOIL isn't used in this problem because if you look closely you're only multiplying -3 by whats in the parenthesis

OpenStudy (anonymous):

so it would be (-3*b) + (-3*4)

OpenStudy (michele_laino):

yes, the application of the distributive property of multiplication over addition is the same as to use foil

OpenStudy (anonymous):

So this is what I'll get\[7b+3-4b=-9-3b = 3b+3=-9-3b\]

OpenStudy (anonymous):

then I go on to get -2 but my teacher said the answer is suppose to be -1/2

OpenStudy (michele_laino):

you should get this: \[3b + 3 = - 3b - 9\] so you are right!

OpenStudy (anonymous):

Thank you, I thought I did it wrong.

OpenStudy (michele_laino):

thank you!